find the area under the standard normal curve between z=0.52 and z=2.31


1
Expert's answer
2021-05-30T22:20:05-0400

"P(0.52<Z<2.31)=P(Z<2.31)-P(Z<0.52)=0.9896-0.6985=0.2911."